Mr ree is 4 years older than ms ree, but their actual ages are a mystery, if they tell you that the summ of the squares of their ages is 3706, how old are they?
Answer by nerdybill(7384)

Let x = Ms Ree's age
then because "Mr ree is 4 years older than ms ree"
x+4 = Mr Ree's age
.
From: "the sum of the squares of their ages is 3706"
x^2 + (x+4)^2 = 3706
x^2 + x^2+8x+16 = 3706
2x^2+8x+16 = 3706
2x^2+8x-3690 = 0
x^2+4x-1845 = 0
.
Factoring:
(x-41)(x+45) = 0
x = {41, -45}
.
We can toss out the negative answer so:
x = 41 (Ms Ree's age)
.
Mr. Ree's age:
x+4 = 41+4 = 45 years (Mr. Ree's age)